# Simultaneous Masking, Not Prompting Optimization: A Paradigm Shift in Fine-tuning LLMs for Simultaneous Translation

### v1

@@ 4.3 Computational Inefficiency
Avoiding KV caching and recomputing all the keys and values at each prediction step is the default solution for resolving the aforementioned fine-tuning-inference mismatch and positional confusion problems while employing prefix fine-tuning. Although effective from a translation quality standpoint, doing so incurs a large computational cost, an undesirable result for streaming tasks like simultaneous translation where latency is equally important.
Outside of KV caching, the computational costs necessary for prefix fine-tuning methods are excessive. By subdividing each sample into multiple, the dataset drastically expands, contributing toward an increased cost to complete each epoch Agostinelli et al. (2023); Koshkin et al. (2024); Wang et al. (2023). Such an increase causes the duration of each epoch to rise by upwards of a factor of 5. However, unlike normal methods of expanding a dataset through data augmentation, prefix fine-tuning does not add additional information. It is from this added computational burden that Agostinelli et al. (2023); Wang et al. (2023) are forced to fine-finetune with a subset of their entire prefix datasets.
Alternatively, methods of restructuring the prompt as in Koshkin et al. (2024); Wang et al. (2024) have computational burdens of their own. For instance, Wang et al. (2024) requires adding delimiting tokens in the prompt sequence, expanding the sequence length. Similarly, the requirement of padding tokens to induce a causal alignment between the source and target sequences, as in Koshkin et al. (2024), also expands the sequence length. Since the computational cost of the self-attention cost in the LLM scales quadratically with the sequence length, such a method is undesirable for both inference and fine-tuning.
Currently, no computationally efficient fine-tuning approach exists that enables computationally efficient inference. Identifying such a method is necessitated by the desire for low latency and high-quality simultaneous translations and reducing the already high computational costs of fine-tuning LLMs.
(a) The attention for the first prediction step.
(b) The inference mirrored attention for the first prediction step.
Figure 2: The matching attention connections during inference and finetuning for simultaneous translation.
@@ 5 SimulMask: A Paradigm Shift
In this work, we propose SimulMask, a paradigm shift in fine-tuning LLMs for simultaneous translation that eschews current methods of prompting optimization. It is through SimulMask, a method for restricting attention connections during fine-tuning, that we efficiently solve the fine-tuning-inference mismatch and positional confusion problem. Although we specifically apply SimulMask for the wait-k decision policy, it is broadly applicable to a range of decision policies. In this section, we will (1) explain the process of restricting attention connections to model simultaneous translation during fine-tuning, (2) describe SimulMask, a masking method for inducing these restricted attention connections, and (3) present a solution for positional confusion that enabled by SimulMask.
@@ 5.1 Inference Mirrored Attention
Under simultaneous translation, the latest translation token at each prediction step is conditioned only on the running source tokens. Specialized attention masks on the Transformer could achieve such conditioning; however, directly mapping these to LLMs is impossible since they fail to enforce autoregressive language modeling and cannot mask properly when the prompt, source, and target sequences are collectively included in the queries and keys. As such, prior works attempted to achieve such conditioning during fine-tuning using prompting optimization strategies littered with shortcomings. We aim to return to a modeling simultaneous translation with attention masks by creating inference mirrored attention. Through inference mirrored attention, we mirror the attention connections during inference at fine-tuning according to the chosen decision policy.
As an example, suppose we model the attention connections for a wait-1 decision policy where the complete oracle input sequence is “$p_{1},s_{1},s_{2},s_{3},s_{4},p_{2},t_{1},t_{2},t_{3},t_{4}$”. In the sequence, $s_{1},s_{2},s_{3},s_{4}$ and $t_{1},t_{2},t_{3},t_{4}$ are the 4-word source and target sequences and $p_{1}$ and $p_{2}$ are prompting regions. Then, at inference, by definition of the wait-1 policy, $p_{2}$ predicts $\hat{t}_{1}$ while conditioned on the partial sequence $p_{1},s_{1},p_{2}$. As such, as shown in Figure 2(a) the query of $p_{2}$ attends to the keys of $p_{1},s_{1},p_{2}$. Thus, during fine-tuning, to eliminate the fine-tuning-inference mismatch, the query of $p_{2}$ is limited to similarly attend to the keys of $p_{1},s_{1},p_{2}$ as shown in Figure 2(b) rather than the entire source sequence. For each successive prediction step, the previously predicted target word, $t_{i}$, predicts the next target word, $t_{i+1}$ by conditioning on an extra source word, $s_{i}$, acquired from the previous read step. To mimic such behavior at fine-tuning, the query for $t_{i}$ attends to identical keys as its inference step. We provide the attention connections for the remainder of the queries in our example in Appendix A.
@@ 5.2 SimulMask
To achieve the above inference mirrored attention, we opt for an attention mask to restrict attention during fine-tuning to mimic an arbitrary decision policy during simultaneous translation. An attention mask is preferable to prompting optimization as it is flexible and directly extends the LLM causal attention mask. We call such an attention mask SimulMask.
As a demonstration, let us create a SimulMask for the wait-1 policy that extends our example from Section 5.1. Since the LLM is autoregressive, SimulMask begins with a causal attention mask from which attention connections are removed to be identical to attention connections during simultaneous translation. If we recall from our example, the prompt, $p_{2}$, generates the first target token, $t_{1}$ conditioned on $p_{1},s_{1},p_{2}$. As such, SimulMask eliminates the attention between $p_{2}$ and $s_{2},s_{3},s_{4}$. Similarly, $t_{1}$ and $t_{2}$ are conditioned on $p_{1},s_{1},s_{2},p_{2},t_{1}$ and $p_{1},s_{1},s_{2},s_{3},p_{2},t_{1},t_{2}$, respectively. Thus, attention connections are eliminated between $t_{1}$ and $s_{3},s_{4}$ and $t_{2}$ and $s_{4}$. The SimulMask for the aforementioned wait-1 policy is provided in Figure 3.
Figure 3: The SimulMask for modeling simultaneous translation according to a wait-1 decision policy during fine-tuning.
Since each decision policy performs read/write decisions differently and each limits attention differently, this requires a unique attention mask for every sentence. However, this can be done straightforwardly. The general procedure to construct a SimulMask for a given policy and sentence consists of the following steps:
@@
The computation for constructing an arbitrary SimulMask is negligible compared with the computation of the LLM’s forward and backward passes during fine-tuning. Since SimulMask is not applied during inference, it does not impact computational cost at deployment. Therefore, SimulMask is a necessary option for mimicking simultaneous translation during fine-tuning and providing low-latency translations at inference.
@@ 5.3 Positional Reordering
${\bm{q}}^{(h)}_{i}{{\bm{K}}^{(h)}}^{T}+{\bm{M}}_{i}+m\cdot[-(i-1),...,-1,0]$ (4)
Though simple, ALiBi has demonstrated an ability to extrapolate to much larger sequence lengths than other state-of-the-art positional encodings, making it desirable for LLMs.
Unfortunately, ALiBi, by default, does not mesh with SimulMask due to SimulMask removing attention connections between the target queries and source keys. These removed attention connections create a gap in ALiBi biases during fine-tuning that are not present at inference. An example of such a gap is provided in Figure 4(a). In Figure 4(a), the gap is present for $q_{4}$ and $q_{5}$.
@@
Figure 4: The ALiBi biases for a 5x5 attention matrix with SimulMask.
To eliminate the bias gap, we modify ALiBi by reducing the bias values of all query rows influenced by SimulMask. For each query row, the reduction in bias values is equivalent to the number of attention connections removed along the row using SimulMask. Figure 4(b) provides an example of such a modification. In the case of $q_{4}$, it is no longer able to attend to $k_{2}$ and $k_{3}$; therefore, the bias on the right of the gap is reduced by 2. In modifying SimulMask, we eliminate positional confusion from the LLM during simultaneous translation.
@@ 6.1 Fine-tuning
Our fine-tuning experiments used the Simul-LLM framework Agostinelli et al. (2023). Each experiment used a 1.3 billion parameter Falcon model pre-trained on the RefinedWeb dataset Penedo et al. (2023). The model architecture consisted of 24 layers. Each self-attention calculation had 32 attention heads, and the hidden size was 2048. We refer to models fine-tuned with prefix fine-tuning that recompute and do not recompute the KV cache as falcon-prefix-rec and falcon-prefix-norec, respectively. The models fine-tuned with SimulMask with and without modifying ALiBi are named as falcon-simul-norec-mod falcon-simul-norec, respectively. Finally, the model fine-tuned with a causal attention mask and evaluated recomputing the KV cache is named as falcon-causal-rec. For comparison purposes, we fine-tuned a Falcon model with a causal attention mask and evaluated it for neural machine translation. We refer to it as falcon-offline.
All fine-tuning was conducted on a single H100 GPU with bfloat16 precision. The non-prefix fine-tuning duration consisted of 2 epochs using a batch size of 64, whereas the prefix fine-tuning duration consisted of 1 epoch using a batch size of 1024. We applied a learning rate of $2e^{-4}$, a weight decay of 0.1, and an inverse square root scheduler. Our optimizer was AdamW Loshchilov and Hutter (2017). The fine-tuned falcon models all used ALiBi. All fine-tuning used the English-French (en-fr), English-Italian (en-it), and English-Dutch (en-nl) language pairs of the IWSLT 2017 training set Cettolo et al. (2017).
@@ 6.2 Evaluation
We evaluated translation quality and latency for simultaneous translation using Simul-LLM inference agents Agostinelli et al. (2023) interfacing with the SimulEval toolkit Ma et al. (2020a). The translation quality was determined using detokenized BLEU with SacreBLEU Post (2018). Alternatively, the latency was determined using Length-Adaptive Average Lagging (LAAL) Papi et al. (2022b). The computational cost of simultaneous translation was recorded with FLOPs, and time in seconds to complete translation. All metrics were obtained on a single A40 GPU with bfloat16 precision. Each model was evaluated at a wait-k four lower, for which it was fine-tuned. For instance, a model fine-tuned for wait-5 was evaluated at wait-1.
@@ 7 Results
Target Length Frequency Pred Time (s) Rec Time (s) Pred GFLOPs Rec GFLOPs
0-10 5.3 % 0.40 0.07 27.60 19.87
10-20 24.5 % 0.80 0.23 46.91 95.08
20-30 24.1 % 1.31 0.44 71.79 252.77
30-40 18.4 % 1.81 0.70 96.54 524.82
40-50 8.6 % 2.39 0.93 124.62 866.00
50-60 7.5 % 2.95 1.21 151.98 1352.39
60-70 3.1 % 3.43 1.41 175.44 1839.24
70-80 2.4 % 3.94 1.42 200.67 1897.82
80-90 2.1 % 4.53 1.67 229.13 2467.18
90-100 1.5 % 5.07 1.84 255.48 2870.21
Table 1: The computation measurements using the wait-3 policy for falcon-causal-rec.
@@ 7.1 Translation Quality and Latency Results
In this section, we demonstrate the efficacy of fine-tuning with SimulMask compared to fine-tuning with a causal attention mask or with prefix fine-tuning using BLEU scores and LAAL. All evaluations are performed across wait-1, wait-3, wait-5, and wait-7 policies. The translation quality and latency results on the English-French, English-Dutch, and English-French language pairs are provided in Figures 5, 6, and 7.
Figure 5: The translation quality plotted against latency for LLMs on the English-French language pair.
Figure 6: The translation quality plotted against latency for LLMs on the English-Dutch language pair.
Figure 7: The translation quality plotted against latency for LLMs on the English-Italian language pair.
Notably, in Figures 5, 6, and 7, falcon-simul-norec-mod outperforms falcon-causal-rec across all wait-k values. On average, the increase across these wait-k values in terms of translation quality is 5.44, 8.38, and 4.78 BLEU on the English-French, English-Dutch, and English-Italian language pairs. Even more impressive, falcon-simul-norec-mod outperforms falcon-prefix-rec at low wait-k values and matches it for higher wait-k values where, on average, across all wait-k values, it increases BLEU score by 7.34, 3.00, and 2.98 BLEU on the English-French, English-Dutch, and English-Italian language pairs. Such increases in translation quality reflect the efficacy of SimulMask in adapting an LLM for simultaneous translation during fine-tuning.
Figures 5, 6, and 7 also provide an ablation demonstrating the importance of modifying ALiBi with SimulMask for high-quality translations by comparing falcon-simul-norec-mod with falcon-simul-norec. For each wait-k value and language pair, falcon-simul-norec-mod outperforms falcon-simul-norec. Quantified when averaged over all wait-k values, falcon-simul-norec-mod provides an increase of 1.20, 4.56, and 3.03 BLEU on the English-French, English-Dutch, and English-Italian language pairs over falcon-simul-norec. Unsurprisingly, at higher wait-k values where the setting approaches neural machine translation, the difference in BLEU scores becomes less pronounced between the models.
A secondary ablation is provided in Figures 5, 6, and 7 by comparing falcon-prefix-rec and falcon-prefix-norec. Doing so demonstrates translation quality increasing by recomputing the KV cache across all wait-k values. When averaged across these wait-k values, the increase is 3.58, 7.48, and 1.72 BLEU on the English-French, English-Dutch, and English-Italian language pairs. Furthermore, as with the previous ablation, the difference in the BLEU score becomes less pronounced for the higher wait-k values.
One observation that initially may seem strange is that models evaluated at lower wait-k values have their LAAL deviate from their respective k to a greater degree than those evaluated at higher wait-k. Such an increase is a byproduct of the lower wait-k models generating longer predictions than their corresponding references. The increased generation length is a byproduct of the model hallucinating on sequences provided insufficient context, causing them to repeat a single token multiple times. The most pronounced occurrence of this hallucination is provided by falcon-prefix-norec and falcon-prefix-rec on the English-French language pair. The output predictions from both these models showcase repeating single tokens or phrases causing not only increased LAAL but decreased BLEU score.
@@ 7.2 Compuational Saving Results
While not using KV caching, the computation for each write step of simultaneous translation comprises of (1) recomputing the keys and values to reflect the source sequence update and (2) generating the new target tokens. In this section, we report the average time and GFLOPs required to recompute keys and values of the predicted target sequence (Rec Time and Rec GFLOPs) and generate the new target tokens (Pred Time and Pred GFLOPs) for each translation. In doing so, we demonstrate the necessity of enabling KV caching for simultaneous translation.
All computational evaluations are performed with falcon-causal-rec using the wait-3 policy on a random sampling of 1000 samples from the English-French language pair of the IWSLT 2017 test set and are reported in Table 1. The computation measurements in Table 1 are an average of samples grouped by the number of tokens in their output translations (Target Length). The frequency that a predicted translation from falcon-causal-rec is of a specified length range is also reported (Frequency). The results for wait-1, wait-5, and wait-7 policies are provided in Appendix B.
From Table 1, in terms of GFLOPs, the recomputation of the keys and values contributes toward a greater portion of the total computation than the new target token generation. However, inversely, more time is spent generating new target tokens than recomputing the keys and values. The reason for the discrepancy is the recomputation of the keys and values is performed in parallel on the A40 GPU whereas the generation of the new target tokens is a sequential operation. On average, across all target lengths, taking into account their frequencies, the recomputation of the KV cache contributed to 26.8% of the computation time and 87.9% of the FLOPs required to generate a target translation at wait-3.

### v4

@@ 4.3 Computational Inefficiency
Avoiding KV caching and instead recomputing all the keys and values at each prediction step is the default solution for resolving the aforementioned fine-tuning-inference mismatch and positional confusion problems while employing prefix fine-tuning. Although effective from a translation quality standpoint, doing so incurs a large computational cost, an undesirable result for streaming tasks like SimulMT, where latency is equally important.
Outside of KV caching, the computational costs necessary for prefix fine-tuning methods are excessive. By subdividing each sample into multiple, the dataset drastically expands, contributing toward an increased cost to complete each epoch Agostinelli et al. (2023); Koshkin et al. (2024); Wang et al. (2023). Such an increase causes the duration of each epoch to rise by upwards of a factor of 10 (as exemplified in Table 1 in Section 7.2). However, unlike normal methods of expanding a dataset through data augmentation, prefix fine-tuning does not actually add additional information. It is from this added computational burden that Agostinelli et al. (2023) and Wang et al. (2023) are forced to fine-tune with a subset of their entire prefix datasets.
Alternatively, methods of restructuring the prompt as in Koshkin et al. (2024) and Wang et al. (2024) have computational burdens of their own. For instance, Wang et al. (2024) requires adding delimiting tokens in the prompt sequence, expanding the sequence length. Similarly, the requirement of padding tokens to induce a causal alignment between the source and target sequences, as in Koshkin et al. (2024), also expands the sequence length. Since the computational cost of the self-attention cost in the LLM scales quadratically with the sequence length, such a method is undesirable for both inference and fine-tuning.
Currently, no computationally efficient fine-tuning approach exists that enables computationally efficient inference. Identifying such a method is necessitated by the desire for low latency and high-quality translations and reducing the already high computational costs of fine-tuning LLMs.
@@ 5 SimulMask: A Paradigm Shift
In this work, we propose SimulMask, which we believe could be a paradigm shift in fine-tuning LLMs for SimulMT that eschews current methods of prompting optimization. By restricting attention during fine-tuning, SimulMask efficiently solves the fine-tuning-inference mismatch and positional confusion problem. We demonstrate SimulMask through its application for the wait-k decision policy, but it should be noted that SimulMask broadly applies to various decision policies.
@@ 5.1 Inference Mirrored Attention
We first introduce the concept of Inference Mirrored Attention that cleverly models SimulMT during LLM fine-tuning. Under SimulMT, the latest translation token at each prediction step is conditioned only on the running source tokens. For conventional Transformers, specialized attention masks could achieve such conditioning; however, directly mapping these to LLMs is impossible since they fail to enforce autoregressive language modeling and cannot mask properly when the prompt, source, and target sequences are collectively included in the queries and keys. As such, prior works attempted to achieve such conditioning during fine-tuning using prompting optimization strategies littered with shortcomings. The proposed inference mirrored attention is aimed to model SimulMT with attention masks for LLMs by mirroring the attention during inference at fine-tuning according to the chosen decision policy.
(a) Inference.
(b) Fine-tuning.
Figure 1: Inference Mirror Attention for matching attention during inference and fine-tuning for SimulMT.
As an example, suppose we model the attention for a wait-1 decision policy where the complete oracle input sequence is “$p_{1},s_{1},s_{2},s_{3},s_{4},p_{2},t_{1},t_{2},t_{3},t_{4}$”. In the sequence, $s_{1},s_{2},s_{3},s_{4}$ and $t_{1},t_{2},t_{3},t_{4}$ are the 4-word source and target sequences and $p_{1}$ and $p_{2}$ are prompting regions. Then, at inference, by definition of the wait-1 policy, $p_{2}$ predicts $t_{1}$ while conditioned on the partial sequence $p_{1},s_{1},p_{2}$. As such, as shown in Figure 1(a) the query of $p_{2}$ attends to the keys of $p_{1},s_{1},p_{2}$. Thus, during fine-tuning, to eliminate the fine-tuning-inference mismatch, the query of $p_{2}$ should be limited to similarly attend to the keys of $p_{1},s_{1},p_{2}$ as shown in Figure 1(b) rather than the entire source sequence. For each successive prediction step, the previously predicted target word, $t_{i}$, predicts the next target word, $t_{i+1}$ by conditioning on an extra source word, $s_{i+1}$, acquired from the previous read step. To mimic such behavior at fine-tuning, the query for $t_{i}$ attends to identical keys as its inference step. The complete steps of this example are in Appendix A.2.
@@ 5.2 SimulMask
To achieve the above inference mirrored attention, we opt for an attention mask to restrict attention during fine-tuning to mimic an arbitrary decision policy during SimulMT. An attention mask is preferable to prompting optimization as it is flexible and directly extends the LLM causal attention mask. We call such an attention mask SimulMask.
As a demonstration, let us create a SimulMask for the wait-1 policy that extends our example from Section 5.1. As depicted in Figure 2, since the LLM is autoregressive, SimulMask begins with a causal attention mask from which attention is limited to be identical to the attention during SimulMT for the source sequence. Starting from the prompt $p_{2}$, from the example in Figure 1(b), $p_{2}$ generates the first target token, $t_{1}$, conditioned on $p_{1},s_{1},p_{2}$. As such, SimulMask eliminates the attention between $p_{2}$ and $s_{2},s_{3},s_{4}$. Similarly, $t_{1}$ and $t_{2}$ are conditioned on $p_{1},s_{1},s_{2},p_{2},t_{1}$ and $p_{1},s_{1},s_{2},s_{3},p_{2},t_{1},t_{2}$, respectively. Thus, attention is eliminated between $t_{1}$ and $s_{3},s_{4}$ and $t_{2}$ and $s_{4}$.
Figure 2: SimulMask for modeling SimulMT according to a wait-1 decision policy during fine-tuning.
SimulMask is a flexible scheme that supports a range of decision policies. Since each decision policy performs read/write decisions differently and each limits attention differently, this requires a unique attention mask for every sentence. However, this can be done straightforwardly. The general procedure to construct a SimulMask for a given policy and sentence consists of the following steps:
@@
As reported in Section 7.2, the computation for constructing an arbitrary SimulMask is minor, and since SimulMask is not applied during inference, it does not impact computational cost at deployment. Therefore, SimulMask is an efficient option for mimicking SimulMT during fine-tuning and providing low-latency translations at inference.
@@ 5.3 Positional Reordering
${\bm{q}}_{i}{\bm{K}}^{T}+{\bm{M}}_{i}+m\cdot[-(i-1),...,-1,0]$ (4)
Though simple, ALiBi has demonstrated an ability to extrapolate to much larger sequence lengths than other state-of-the-art positional encodings, making it desirable for LLMs like Falcon Penedo et al. (2023), BLOOM Le Scao et al. (2023), etc.
Unfortunately, ALiBi, by default, does not mesh with SimulMask as SimulMask removes attention between the target queries and source keys. This removed attention creates a gap in ALiBi biases during fine-tuning that are not present at inference. An example of such a gap is provided in Figure 3(a), where both $q_{4}$ and $q_{5}$ have gaps in the position distance.
@@
Figure 3: ALiBi biases with SimulMask.
To eliminate the bias gap, we modify ALiBi by reducing the bias values of all query rows influenced by SimulMask. For each query row, the reduction in bias values is equivalent to the amount of attention removed along the row using SimulMask. Figure 3(b) provides an example of such a modification. In the case of $q_{4}$, it is no longer able to attend to $k_{2}$ and $k_{3}$; therefore, the bias on the right of the gap is reduced by 2. Together with the modified ALiBi, SimulMask eliminates positional confusion from the LLM during SimulMT.
@@ 6 Experimental Setup
(a) English-French language pair results.
(b) English-Dutch language pair results.
(c) English-Italian language pair results.
(d) English-Romanian language pair results.
(e) English-German language pair results.
Figure 4: Translation quality plotted against latency for LLMs on the English-French, English-Dutch, English-Romanian, and English-German language pairs.
@@ 6.1 Fine-tuning
Our fine-tuning was conducted with the Simul-LLM framework Agostinelli et al. (2023), which contains our publicly available code for SimulMask [^1] https://github.com/OSU-STARLAB/Simul-LLM. Each experiment used a 1.3 billion parameter Falcon model pre-trained on the RefinedWeb dataset Penedo et al. (2023). We compared 7 schemes:
  • causal-offline: Fine-tuned with a causal attention mask and evaluated for NMT (non-SimulMT).
  • causal-rec: Fined-tuned with a causal attention mask and evaluated with recomputing the KV cache.
  • prefix-rec and prefix-norec: Fined-tuned with prefix fine-tuning and evaluated with/without recomputing the KV cache.
  • converse-norec: Fined-tuned with a conversational prompting strategy and evaluated without recomputing the KV cache.
  • SM-norec-mod and SM-norec: Fined-tuned with SimulMask with/without modifying ALiBi and evaluated without recomputing the KV cache.
Appendix A.3 provides all model hyperparameters. Our fine-tuning experiments included the English-French (en-fr), English-Italian (en-it), English-Dutch (en-nl), English-Romanian (en-ro), and English-German (en-de) language pairs of the IWSLT 2017 dataset Cettolo et al. (2017).
@@ 6.2 Evaluation
We evaluated translation quality and latency for SimulMT using Simul-LLM inference agents Agostinelli et al. (2023) interfacing with the SimulEval toolkit Ma et al. (2020a). The translation quality was determined using detokenized BLEU with SacreBLEU Post (2018) and chrF++ Popović (2017). Latency was determined using Length-Adaptive Average Lagging (LAAL) Papi et al. (2022b). The computational cost of SimulMT was recorded with GFLOPs. All metrics were obtained on a single A10 GPU with bfloat16 precision. The models fine-tuned for the wait-k policy were evaluated at a wait-k four lower, for which they were fine-tuned, as suggested by Ma et al. (2019).
@@ 7.1 Translation Quality and Latency Results
In this section, we demonstrate the efficacy of fine-tuning with the proposed SimulMask compared with other schemes using BLEU scores and LAAL. All wait-k model evaluations are performed across wait-{1,3,5,7}, and the converse-norec is evaluated for a chunk size of 1, 3, 5, 7, 9, and 11. Figure 4 provides the BLEU translation quality and latency results on the English-French, English-Dutch, English-Italian, English-Romanian, and English-German language pairs. We provide the numerical BLEU and chrF++ translation quality results in Tables 3 and 2 of Appendix A.4.
Overall, throughout Figure 4, the proposed SM-norec-mod outperforms or matches the translation quality of causal-rec, prefix-rec, and converse-norec across all latencies. The only major exception occurs at wait-1, where converse-norec outperforms SM-norec-mod on the English-Romanian language pair. This overall excellent performance in terms of translation quality underscores the importance of the proposed method.
Furthermore, Figure 4 provides two ablation studies. The first ablation demonstrates the importance of modifying ALiBi with SimulMask for high-quality translations by comparing SM-norec-mod with SM-norec. For each wait-k value and language pair, SM-norec-mod outperforms SM-norec. Unsurprisingly, at higher wait-k values where the setting approaches NMT, the difference in BLEU scores becomes less pronounced between the models.
A secondary ablation is provided in Figure 4 by comparing prefix-rec and prefix-norec. Doing so demonstrates that translation quality increases by recomputing the KV cache across all wait-k values. Similarly, as with the previous ablation, the difference in the BLEU score becomes less pronounced for the higher wait-k values.
An interesting observation is that models evaluated at lower wait-k values have their LAAL deviate from their respective k to a greater degree than those evaluated at higher wait-k. Such an increase is a byproduct of the lower wait-k models generating longer predictions than their corresponding references. The increased generation length is a byproduct of the model hallucinating on sequences provided insufficient contexts. These hallucinations are most noticeable with prefix-rec and prefix-norec in Figure 4(e).
@@ 7.2 Compuational Saving Results
Fine-tuning LLMs with SimulMask also features reduced training time compared with LLMs leveraging prefix fine-tuning or conversational prompting. For instance, this is reflected in the fine-tuning times for one epoch on an H100 GPU on the English-French dataset of the IWSLT 2017 dataset, as reported in Table 1 Cettolo et al. (2017).
Fine-tuning Approach Time (s)
Prefix Fine-tuning 9953
Conversational Prompting 1274
SimulMask 1014
Causal Mask 727
Table 1: Time to complete one epoch for different fine-tuning approaches on an H100.
Furthermore, we find that SM-norec is also more computationally efficient at inference than prefix-rec and converse-norec. We report these results in GFLOPs that are needed to complete a sentence translation in Figure 5. The data used to obtain the results was a random 1000 samples from the English-French split of the IWSLT 2017 test set Cettolo et al. (2017). The models chosen either used wait-3 or a chunk size of 5.
005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 5002,00020002{,}0002 , 0002,50025002{,}5002 , 500SM-norec-modconverse-norecprefix-reccausal-recCompuation (GFLOPs)
Figure 5: Box plots of the computational cost of each method in GFLOPs during inference.
By leveraging SimulMask during fine-tuning, we eschew the need to recompute the KV cache at inference. In doing so, SimulMask saves computation compared to prefix-rec and causal-rec. We demonstrate the proportions of computation in GFLOPs dedicated to re-computing the KV cache and processing/predicting initial tokens in Figure 6 (based on prefix-rec). The sequence length is the number of tokens in the predicted target and input source. As can be seen, it is critical to avoid recomputing KV cache, as achieved by SimulMask, to provide low latency translations, especially at longer sequence lengths.
0050505050100100100100002,00020002{,}0002 , $0004{,}$  [^0] Sequence LengthComputation (GFLOPs)InitialRecompute
Figure 6: Separated computational cost in GFLOPs between initial (or required) computational cost and the cost of recomputing already emitted target words in a provided prompt during translation versus the sequence length of a given sample.